OpenStudy (anonymous):

Solve the system by the substitution method. xy = 12 x2 + y2 = 40 A. {( 2, 6), ( 6, 2), ( 2, -6), ( 6, -2)} B. {( 2, 6), ( -2, -6), ( 2, -6), ( -2, 6)} C. {( 2, 6), ( -2, -6), ( 6, 2), ( -6, -2)} D. {( -2, -6), ( -6, -2), ( -2, 6), ( -6, 2)} What substitution method?

OpenStudy (anonymous):

set the second equation equal to x and plug in the answer to the x in the first equation

OpenStudy (anonymous):

The method of solving "by substitution" works by solving one of the equations for one of the variables, and then plugging this back into the other equation, "substituting" for the chosen variable and solving for the other. Then you back-solve for the first variable.

OpenStudy (anonymous):

\[xy = 12\] \[x^2 + y^2 = 40\] so I could get \[y^2 = 40 - x^2\] ?

OpenStudy (anonymous):

then I have to lose the sq on y before plugging it in for y right?

OpenStudy (anonymous):

so x=20-y and plug that into the first equation so (20-y)y=12 then solve that so 20y-y^2=12 and oolve

OpenStudy (anonymous):

am I supposed to have a number that plugs in to those and equals 12?

OpenStudy (anonymous):

oh Is that squared? I thought it was 2x and 2y

OpenStudy (anonymous):

it's sq, sorry... I corrected it in a comment above